### v3

# LiveScene: Language Embedding Interactive Radiance Fields for Physical Scene Rendering and Control

Abstract
This paper scales object-level reconstruction to complex scenes, advancing interactive scene reconstruction. We introduce two datasets, OmniSim and InterReal, featuring 28 scenes with multiple interactive objects. To tackle the challenge of inaccurate interactive motion recovery in complex scenes, we propose LiveScene, a scene-level language-embedded interactive radiance field that efficiently reconstructs and controls multiple objects. By decomposing the interactive scene into local deformable fields, LiveScene enables separate reconstruction of individual object motions, reducing memory consumption. Additionally, our interaction-aware language embedding localizes individual interactive objects, allowing for arbitrary control using natural language. Our approach demonstrates significant superiority in novel view synthesis, interactive scene control, and language grounding performance through extensive experiments. Project page: https://livescenes.github.io.
Abstract
This supplementary material accompanies the main paper by providing more details for reproducibility as well as additional evaluations and qualitative results to to verify the effectiveness and robustness of LiveScene:
$\triangleright$ Section. 7: Configurations of OmniSim and InterReal dataset, including scene assets, interaction variables generation, mask and prompts annotation, and dataset visualization.
$\triangleright$ Section. 8: Video demonstration and anonymous link: https://livescenes. github.io.
$\triangleright$ Section. 9: Additional implementation details.
$\triangleright$ Section. 10: Additional experimental results, including more ablation studies, detailed view synthesis quality comparison, interactive scenes geometry comparison and language grounding comparison.
Figure 1: LiveScene enables scene-level reconstruction and control with language grounding. Left: Language-interactive articulated object control in Nerfstudio. Right: LiveScene achieves SOTA rendering quality on OmniSim dataset and exhibits a significant advantage in parameter efficiency.
1 Introduction
Interactive objects are prevalent in our daily lives, and modeling interactable scenes from the real physical world plays an essential role in various research fields, including content generation [34, 37, 55], animation [61, 39, 31], virtual reality [53, 8, 33], robotics [47, 73, 74, 35, 46], and world understanding [20, 14, 56, 19, 11]. This paper tackles the challenging and rarely explored task of reconstructing and controlling multiple interactive objects in complex scenes from a single, casually captured monocular video without previous independent modeling of geometry and kinematics. Prior research on interactable scene modeling, such as CoNeRF [21] and K-Planes [9], typically adopts a joint modeling approach, combining spatial coordinates and all interaction variables as input and representing interactive scene by either implicit MLPs or feature planes. Meanwhile, CoGS [65] learns parameter offsets for different scene parts using multiple independent MLPs after establishing a 3D deformable Gaussian scene. However, these methods primarily focus on capturing interactions for a single object within a clear background, such as a single drawer, toy car, or face [68, 21, 65, 70]. As modeling extends from single objects to multiple objects in complex scenes, as shown in Figure. 1, the interaction spaces become increasingly high-dimensional, complicating these methods for accurate modeling and significantly increasing computational time and memory cost, e.g., 4$\times$ A100 GPU for 2 weeks to converge training in CoNeRF [21] and 500M Gaussian storage for a regular indoor living room in CoGS [65]. Moreover, natural language is an intuitive and necessary interface for interacting with 3D scenes, but language embedding of interactive scenes faces an even more daunting challenge: interaction variation inconsistency. For instance, methods like LERF [24], and OpenNeRF [69], which distill CLIP features into static 3D fields, suffer from significant failures when confronted with scene topology structure changes induced by interactions, such as the distinct structures variation of a cabinet before and after opening.
To address these challenges, we propose LiveScene, the first scene-level language-embedded radiance fields, which compresses high-dimensional interaction spaces into compact 4D feature planes, reducing model parameters while improving optimization effectiveness. LiveScene models multiple object interactions via novel high-dimensional factorization, decomposing the scene into local deformable fields that model individual objects with multi-scale 4D deformable feature planes. To achieve independent control, we introduce a multi-scale interaction probability sampling strategy for factorized local deformable fields. Our interaction-aware language embedding method generates varying language embeddings to localize and control objects under arbitrary states, enabling natural language control. Finally, we construct the first scene-level physical interaction datasets, OmniSim and InterReal, featuring 28 scenes with 70 interactive objects for evaluation.
Experiment results show that our approach achieves SOTA novel view synthesis quality, outperforming existing best methods by +9.89, +1.30, and +1.99 in PSNR on the CoNeRF Synthetic, OmniSim #chanllenging, and InterReal #chanllenging subsets, respectively. Surpassing LeRF [24], LiveScene significantly improves language grounding accuracy by +65.12 of mIOU on the OmniSim dataset. Notably, our method maintains a lightweight, constant model parameter of 39M, scaling well with increasing scene complexity, as shown in Figure. 1. Contributions can be summarized as:
  • We propose LiveScene, the first scene-level language-embedded interactive radiance field, which efficiently reconstructs and controls complex physical scenes, enabling manipulation of multiple articulated objects and language-based interaction.
  • We propose a factorization technique that decomposes interactive scenes into local deformable fields and samples relevant 3D points, enabling control of individual objects. Additionally, we introduce an interaction-aware language embedding method that generates varying embeddings, allowing for language-based control and localization.
  • We construct the first scene-level physical interaction dataset OmniSim and InterReal, containing 28 subsets and 70 interactive objects for evaluation. Extensive experiments demonstrate our SOTA performance and robust interaction capabilities.
2 Related Work
Dynamic Scene Representation. Extending NeRF [40] to dynamic scene reconstruction has made significant progress. Related methods can generally be categorized into time-varying methods, deformable-canonical methods, and hybrid representation methods. The time-varying methods [6, 43, 57, 67] typically model the radiance field directly over time, but struggle to separate dynamic and static objects. Deformable-canonical methods [10, 32, 44, 63] decouple dynamic deformable field and static canonical space, modeling 4D by warping points with deformable field to query the canonical features. However, these methods face challenges in scene topology changes [9]. Hybrid representation methods, on the other hand, have achieved high-quality reconstruction and fast rendering by utilizing time-space feature planes [50, 9, 2], 4D hash encoding [59], dynamic voxels [60], or triple fields [52]. Recently, several works [38, 64, 62, 28] have introduced 3D gaussians [22] into dynamic scene reconstruction, achieving high-quality real-time rendering speeds. However, these methods are limited to reconstructing dynamic scenes and lack the ability to control and understand interactive scenes.
3D Vision-language Fields. Vision-language foundational models [48, 42, 26] with strong generalizability and adaptability inspires numerous language embedded scene representation for 3D scene understanding [1, 72], such as open-vocabulary segmentation [36, 13, 58, 25], 3D visual question answering [15, 5, 18, 17], and 3D language grounding [16, 49, 4, 17]. LeRF [24] is the first to achieve open-vocabulary 3D queries by combining CLIP [48] and DINO [3] with NeRF through feature distillation. Open-NeRF [69] introduces an integrate-and-distill paradigm and leverages hierarchical embeddings to address 2D knowledge-distilling issues from SAM. LEGaussians [51] and LangSplat [45] integrate semantic features into 3D gaussians [22] and achieve precision language query and efficient rendering. However, these methods are limited to static scene understanding and fail to generalize when the interactive scene topology changes.
Controllable Scene Representation. Manipulating reconstructed assets or neural fields is of significant importance for avatar and robotic tasks [7, 12, 21, 29]. CoNeRF [21] pioneered this effort by extending HyperNeRF [44] and introduce a fine-grained controlable neural field with 2D attribute mask and value annotations. CoNFies [66] proposes an automatic controllable avatar system and accelerates rendering by distilling. More recently, CoGS [65] leveraged 3D Gaussians [22] to achieve real-time control of dynamic scenes without requiring explicit control signals. However, these methods typically lack natural language interaction capabilities, relying solely on manual control. Furthermore, most works focus on single or few object interactions, disregarding the interaction between different scene parts, limiting their real-world applications.
3 Methodology
Figure 2: The overview of LiveScene. Given a camera view and control variable $\boldsymbol{\kappa}$ of one specific interactive object, a series of 3D points are sampled in a local deformable field that models the interactive motions of this specific interactive object, and then the interactive object with novel interactive motion state is generated via volume-rendering. Moreover, an interaction-aware language embedding is utilized to localize and control individual interactive objects using natural language.
We aim to establish a representation that models $\alpha$ interactive articulated objects in a complex scene from a monocular video via a rendering-based self-supervised manner. Control variables $\boldsymbol{\kappa}=\left[\kappa_{1},\kappa_{2},...,\kappa_{\alpha}\right]$ indicating object motion states and camera poses of each video frame are given. The overview of LiveScene is shown in Figure. 2. Section. 3.1 introduces the high-dimensional interactive space modeling and challenges. Section. 3.2 presents a multi-scale interactive space factorization and sampling strategy to compress the high-dimensional interactive space into local 4D deformable fields, and model complicated interactive motions of individual objects. Section. 3.3 introduces an interaction-aware language embedding method to localize and control interactive objects using natural language.
3.1 Interactive Space
Assuming a non-rigidly interactive scene with $\alpha$ control variables $\boldsymbol{\kappa}=\left[\kappa_{1},\kappa_{2},...,\kappa_{\alpha}\right]$ corresponding to $\alpha$ objects, we delineate its representation by a high-dimensional function:
$\mathbf{y}=\rho(\mathbf{x},\boldsymbol{\kappa},\mathcal{H};\boldsymbol{\theta}),$ (1)
where $\rho$ is the model of the representation, $\mathbf{x}\in\mathbb{R}^{3}$ are spatial coordinates, $\boldsymbol{\kappa}\in\mathbb{R}^{\alpha}$ are control variables, $\mathcal{H}$ is a set of optional additional parameters (e.g., the view direction), and $\boldsymbol{\theta}$ stores the scene information. The function outputs scene properties $\mathbf{y}$ for the given position $\mathbf{x}$ and $\boldsymbol{\kappa}$ sampling from a ray $\mathbf{r}$, where $\mathbf{y}$ can be represented with color, occupancy, signed distance, density, and BRDF parameters. This paper focuses on color, probability, and language embedding. Distinguishing from 3d static scene or 4d dynamic scene modeling, the sampling point $\mathbf{p}=\left[\mathbf{x}|\boldsymbol{\kappa}\right]\in\mathbb{R}^{3+\alpha}$ in the interactive scene is high-dimensional and variable in topological structure, complicating the scene feature storage and the optimization of representation model, leading to significant time-consuming or memory-intensive training in [21, 65].
3.2 Multi-scale Interaction Space Factorization
For the $(3+\alpha)$-dimensional interactive space containing $\alpha$ control variables, we aim to explicitly represent the high-dimensional space in a concise and compact storage, thereby reducing memory usage and improving optimization. As illustrated in Figure. 2 and Figure. 10, objects exhibit mutual independence, and interaction features are distributed in the $(3+\alpha)$-dimensional interactive space and aggregate into cluster centers. Thus, there exists a set of hyperplanes that partition the space into disjoint regions, with each region containing a local 4D deformable field, as shown in Figure. 3. Hence, the interaction features at sampling point $\mathbf{p}\in\mathbb{R}^{3+\alpha}$ can be projected into a compact 4-dimensional space $\mathbb{R}^{4}$ in Figure. 3 by a transformation.
Figure 3: Illustration of hyperplanar factorization for compact storage. We maintain multiple local deformable fields for each interactive object region $\mathcal{R}_{i}$, and project high-dimensional interaction features into a compact 4D space, which can be further compressed into multiscale feature planes.
Multi-scale Interactive Ray Sampling. We consider using ray sampling to perform the projection transformation. As shown in Figure. 3, assuming a ray $\mathbf{r}(t)=\left[\mathbf{o_{x}}|\mathbf{o}_{\boldsymbol{\kappa}}\right]+t\left[\mathbf{d_{x}}|\mathbf{d}_{\boldsymbol{\kappa}}\right]$ with origin $\mathbf{o_{x}}\in\mathbb{R}^{3},\mathbf{o}_{\boldsymbol{\kappa}}\in\mathbb{R}^{\alpha}$, and direction $\mathbf{d_{x}}\in\mathbb{R}^{3},\mathbf{d}_{\boldsymbol{\kappa}}\in\mathbb{R}^{\alpha}$, the ray intersects with the interaction region at a point $\mathbf{p}=\left[\mathbf{x}|\boldsymbol{\kappa}\right]\in\mathbb{R}^{3+\alpha}$, where $\mathbf{x}\in\mathbb{R}^{3}$ is the 3D position and $\boldsymbol{\kappa}\in\mathbb{R}^{\alpha}$ is the interaction variables. For a given intersection point $\mathbf{p}$, the deformable features can be retrieved from the corresponding local 4D deformable field by maximizing sampling probability $\mathbf{P}$:
$\mathbf{p}_{u}=\left[\mathbf{x}|\boldsymbol{\kappa}(u)\right],\quad u=\operatorname*{arg\,max}_{i}\{\mathbf{P}_{i}\},\quad\mathbf{P}=\Theta(\boldsymbol{\kappa},\boldsymbol{\theta}),$ (2)
where $\mathbf{p}_{u}$ and $\boldsymbol{\theta}$ are the 4D sampling point and probability features at position $\mathbf{x}$ from 3D feature planes. The maximum argument operation of probability decoder $\Theta(\boldsymbol{\kappa},\boldsymbol{\theta})$ maps the interaction variables $\boldsymbol{\kappa}$ to the most probable cluster region in the 4D space, and can be optimized by minimizing the focal loss $\mathcal{L}_{\text{focus}}$ of mask across all the training camera views:
$\mathcal{L}_{\text{focus}}=\beta\cdot\left(1-e^{\sum_{i=1}^{\alpha}\mathbf{M}_{i}\log(\hat{\mathbf{P}}_{i})}\right)^{\gamma}\cdot\left(-\sum_{i=1}^{\alpha}\mathbf{M}_{i}\log(\hat{\mathbf{P}}_{i})\right),$ (3)
where $\mathbf{M}$ is the ground truth mask label, $\hat{\mathbf{P}}$ is the probability map rendering from the interactive probability field, $\beta$ is the balancing factor, and $\gamma$ is the focusing parameter.
Next, the deformable features are used to render local deformable scene color at the sampling point $\mathbf{p}$. In this way, the high-dimensional interaction features are factorized into a 4D space by a lightweight transformation modeling with interaction probability decoder and 3D feature planes in Figure. 2, supervised by deformable masks $\mathbf{M}$. Moreover, leveraging K-Planes [9], the multiple 4D local deformable space can be further compressed in only $\mathbf{C}_{4}^{2}=6$ feature planes. We iteratively sample from coarse to fine within the multi-scale feature plane, retrieving the maximum probability interaction variables $\boldsymbol{\kappa}_{u}$ and indices $u$ at each scale.
Feature Repulsion and Probability Rejection. A latent challenge is that optimizing the interaction probability decoder with varying masks can lead to blurred boundaries in the local deformable field, further causing ray sampling and feature storage conflicts. As illustrated in Figure. 4(a), consider two adjacent local deformable regions $\mathcal{R}_{i}$ and $\mathcal{R}_{j}$, and a point $\mathbf{p}$ in high-dimensional space, suppose $\mathbf{p}$ moves from the cluster center of $\mathcal{R}_{i}$ towards the cluster center of $\mathcal{R}_{j}$, then the probability of $\mathbf{p}$ belonging to $\mathcal{R}_{i}$ gradually decreases, while the probability of $\mathbf{p}$ belonging to $\mathcal{R}_{j}$ increases. To avoid sampling conflicts and feature oscillations at the boundaries, we introduce a repulsion loss for ray pairs $(\mathbf{r}_{i},\mathbf{r}_{j})$, and amplify the feature differences between distinct deformable regions, promoting the separation of deformable field:
$\mathcal{L}_{\text{repuls}}=\mathbf{ELU}(K-\left\|(\mathbf{M}_{i}\odot\mathbf{M}_{j})(\mathcal{F}_{i}-\mathcal{F}_{j})\right\|),\end{aligned}$
where $K$ is a constant hyperparameter, $\mathbf{M}_{i}$ and $\mathbf{M}_{j}$ are the ground truth mask of rays. $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ are the last-layer features of interaction probability decoder in Figure. 2. During training, we randomly select ray pairs and apply $\mathcal{L}_{\text{repuls}}$ to enforce the separation of interactive probability features across local deformable spaces. The initiative is inspired by [25], which demonstrated the efficacy of repulsive forces in disambiguating 3D segmentation results.
Additionally, the probability rejection is proposed to truncate the low-probability samples if the maximum deformable probability $\mathbf{P}$ at sample $\mathbf{p}$ is smaller than threshold $s$, and selects the background feature directly. The operation is defined as:
$u=\left\{\begin{array}[]{ll}\operatorname*{arg\,max}_{i}\{\mathbf{P}_{i}\},&\text{if}\quad\mathbf{P}_{i}\geq s\\ -1,&\text{ otherwise }\end{array}\right..$ (5)
As shown in Figure. 4(b), the proposed operations help the model achieve higher rendering quality, demonstrating their effectiveness in alleviating boundary sampling conflicts.
Figure 4: Illustration of a) boundary sampling conflicts, b) rendering quality comparison.
3.3 Interaction-Aware Language Embedding
Language embedding in interactive scenes is complex and storage-intensive, as 3D distillation faces the dual challenges of high-dimensional optimization and interaction scene variation inconsistency, such as the distinct topological structures of a transformer toy before and after transformation, leading to the failure of SAM [26] segmentation or LERF [24] grounding. As shown in Figure. 2, leveraging the proposed multi-scale interaction space factorization of 3.2, we efficiently store language features in lightweight planes by indexing them according to maximum probability sampling instead of 3D fields in LERF. For any sampling point $\mathbf{p}$, we project it onto $\mathbf{p}_{u}=\left[\mathbf{x}|\boldsymbol{\kappa}(u)\right]$, retrieving a local language feature group by index $d$, and perform bilinear interpolation using $\boldsymbol{\kappa}_{u}$ to obtain a language embedding that adapts to interactive variable changes from surrounding clip features. By interpolating language embeddings, our method not only perceives topological structure changes but also achieves a storage complexity of $\mathbf{O}(C\times\alpha\times\mathbf{dim})$, much smaller than language distillation methods like LERF that operate in 3D scenes, where $\mathbf{dim}$ is the dimension of CLIP feature.
4 Dataset
Figure 5: Overview of the OmniSim and InterReal datasets.
To our knowledge, existing view synthetic datasets for interactive scene rendering are primarily limited to a few interactive objects [68, 21, 65, 70] due to necessitating a substantial amount of manual annotation of object masks and states, making it impractical to scale up to real scenarios involving multi-object interactions. To bridge this gap, we construct two scene-level, high-quality annotated datasets to advance research progress in reconstructing and understanding interactive scenes: OmniSim and InterReal, as shown in Figure. 5. Besides, we use the CoNeRF Synthetic and Controllable [21] dataset for evaluation as well. 1) OmniSim Dataset is rendered through OmniGibson [30] simulator, leveraging 7 indoor scene models: #rs, #ihlen, #beechwod, #merom, #pomaria, #wainscott and #benevolence. By varying the rotation vectors of the articulated objects’ joints and the camera’s trajectory within the scene, we generated 20 high-definition subsets, each consisting of RGBD images, camera trajectory, interactive object masks, and corresponding object state quantities at each time step. 2) InterReal Dataset is captured from 8 real Interactable scenes and finely annotated with interaction variables and masks, camera poses encompassing multiple objects, and articulated motion variables. More details can be found in the supplementary.
5 Experiment
Baselines. We compare LiveScene with the existing 3D static rendering methods [40, 41, 23, 22], 4D deformable methods [9, 44, 43], and controllable scene reconstruction methods [21, 65]. Note that we reimplemented CoGS [65] based on Deformable Gaussian [64] since the official code is unavailable. Additionally, we extended K-Planes [9] from $\mathbf{C}_{4}^{2}$ planes to $\mathbf{C}_{3+\alpha}^{2}$ planes, denoted as MK-Planes, where $\alpha$ represents the number of interactable objects in dynamic scenes. By leveraging the fact that each instance occupies a distinct region, we further compressed the model, denoted as MK-Planes, requiring only $3+3\alpha$ planes.
Implementation. LiveScene is implemented in Nerfstudio [54] from scratch. We represent the field as a multi-scale feature plane with resolutions of $512\times 256\times 128$, and feature dimension of 32. The proposal network adopts a coarse-to-fine sampling process, where each sampling step concatenates the position feature and the state quantity as the query for the 4D deformation probability field, which is a 1-layer MLP with 64 neurons and ReLU activation. We use the Adam optimizer with initial learning rates of 0.01 and a cosine decay scheduler with 512 warmup steps for all networks. The model is trained with an NVIDIA A100 GPU for 80k steps on the OmniSim dataset and 100k steps on the InterReal dataset, using a batch size of 4096 rays with 64 samples. More implementation can be found in the supplemental materials.
5.1 View Synthesis Quality Comparison
Evaluation on CoNeRF Synthetic and Controllable Datasets. We report the quantitative results on CoNeRF Synthetic and Controllable scenes in Table. 1. LiveScene outperforms all the existing PSNR, SSIM, and LPIPS metrics methods on CoNeRF Synthetic scenes with a large margin. In particular, LiveScene achieves 43.349, 0.986, and 0.011 in PSNR, SSIM, and LPIPS, respectively, outperforming the second-best method by 9.894, 0.009, and 0.053. On CoNeRF Controllable, LiveScene achieves the best PSNR of 32.782 and comparable SSIM and LPIPS to the SOTA methods. According to the novel view synthesis results in Figure. 6, LiveScene achieves more detailed and higher rendering quality, demonstrating the effectiveness of LiveScene in modeling object-level interactive scenarios.
Table 1: Quantitative results on CoNeRF synthetic and controllable datasets. LiveScene achieves the best results in all metrics on synthetic scenes and the best PSNR on the controllable datasets.
Method CoNeRF Syntetic CoNeRF Controllable
[…]
NeRF [40] 25.299 0.843 0.197 28.795 0.951 0.210
InstantNGP [41] 27.057 0.903 0.230 26.391 0.884 0.278
3DGS [22] 32.576 0.977 0.077 25.945 0.834 0.414
NeRF + Latent [40] 28.447 0.939 0.115 32.653 0.981 0.182
Nerfies[43] 32.274 0.981 0.180
HyperNeRF[44] 25.963 0.854 0.158 32.520 0.981 0.169
K-Planes [9] 33.301 0.933 0.150 31.811 0.912 0.262
$-\mathcal{M}$[21] 27.868 0.898 0.155 32.061 0.979 0.167
CoNeRF[21] 32.394 0.972 0.139 32.342 0.981 0.168
CoGS [65] 33.455 0.960 0.064 32.601 0.983 0.164
LiveScene (Ours) 43.349 0.986 0.011 32.782 0.932 0.186
Table 2: Quantitative results on OmniSim Dataset. LiveScene outperforms prior works on most metrics and achieves the best PSNR on the #challenging subset with a significant margin.
Method #Easy Sets #Medium Sets #Challenging Sets #Avg (all 20 Sets)
[…]
NeRF [40] 25.817 0.906 0.167 25.645 0.928 0.138 26.364 0.927 0.128 25.776 0.916 0.153
InstantNGP [41] 25.704 0.902 0.183 25.627 0.930 0.140 26.367 0.920 0.143 25.706 0.914 0.164
HyperNeRF [44] 30.708 0.908 0.316 31.621 0.936 0.265 27.533 0.897 0.318 30.748 0.917 0.299
K-Planes [9] 32.841 0.952 0.093 32.548 0.954 0.100 29.833 0.937 0.118 32.573 0.952 0.097
CoNeRF [21] 32.104 0.932 0.254 33.256 0.951 0.207 30.349 0.923 0.238 32.477 0.939 0.234
MK-Planes 31.630 0.948 0.098 31.880 0.951 0.104 26.565 0.887 0.218 31.477 0.946 0.106
MK-Planes 31.677 0.948 0.098 32.165 0.952 0.099 29.254 0.933 0.119 31.751 0.949 0.099
CoGS [65] 32.315 0.961 0.108 32.447 0.965 0.086 28.701 0.970 0.073 32.187 0.963 0.097
LiveScene (Ours) 33.221 0.962 0.072 33.262 0.965 0.072 31.645 0.948 0.093 33.158 0.962 0.074
Table 3: Quantitative results on InterReal Dataset. Our method outperforms others in most settings, with a significant advantage of PSNR, SSIM, and LPIPS on the #challenging subset.
Method #Medium Sets #Challenging Sets #Avg (all 8 Sets)
[…]
NeRF [40] 20.816 0.682 0.190 21.169 0.728 0.337 20.905 0.694 0.227
InstantNGP [41] 21.700 0.776 0.215 21.643 0.745 0.338 21.686 0.769 0.245
HyperNeRF [44] 25.283 0.671 0.467 25.261 0.713 0.517 25.277 0.682 0.480
K-Planes [9] 27.999 0.813 0.177 26.427 0.756 0.331 27.606 0.799 0.215
CoNeRF [21] 27.501 0.745 0.367 26.447 0.734 0.472 27.237 0.742 0.393
CoGS [65] 30.774 0.913 0.100 30.774 0.913 0.100
LiveScene (Ours) 30.815 0.911 0.066 28.436 0.846 0.185 30.220 0.895 0.096
Evaluation on OmniSim Datasets. OmniSim dataset is categorized into 3 interaction level subsets: #easy, #medium, and #challenging, based on the number of interactive objects in each scene. As shown in Table. 2, LiveScene achieves the best PSNR, SSIM, and LPIPS on all interaction level subsets of OmniSim, with average PSNR, SSIM, and LPIPS of 33.158, 0.962, and 0.074, respectively. Notably, substantial performance degradation is observed across all methods as the quantity and complexity of interactive objects increase, e.g., CoGS [65] experiences a 3.641 dB PSNR drop from #easy to #challenging. While LiveScene maintains a relatively stable high performance across all subsets, demonstrating its robustness in modeling complex interactive scenarios.
Evaluation on InterReal Datasets. We divided InterReal dataset into #medium and #challenging subsets. In  Table. 3, CoGS [65] underperforms compared to LiveScene on the #medium subset and fails to converge when faced with long camera trajectories and a large number of interactive objects in the scene (#challenging), highlighting the limitation of existing methods in modeling real-world interactive scenarios. In contrast, LiveScene achieves the highest PSNR of 28.436 and the lowest LPIPS of 0.185 on the #challenging subset, indicating its superiority in modeling real-world large-scale interactive scenarios.
Figure 6: View Synthesis Visualization on CoNeRF Controllable Dataset. The proposed method achieves higher-quality rendering results compared with the existing methods.
Figure 7: View Synthesis Visualization on OmniSim Dataset. We compare our method with SOTA methods on RGB rendering across three scenes: #rs, #ihlen, and #pomaria. Boxes of different colors represent distinct interactive objects within the scene.
View Synthesis Visulization. Figure. 7 presents the novel view synthesis results of LiveScene and the SOTA methods on OmniSim dataset. The results reveal that LiveScene generates more detailed results than SOTA methods, particularly in complex interactive scenarios. For instance, on the #pomaria scene featuring an openable dishwasher, CoNeRF [21] fails to capture details, while CoGS [65] and MK-Planes exhibit residual artifacts. In contrast, our method accurately reconstructs the internal details. Another challenge arises in the #rs scene, where other methods struggle to reconstruct distant and static objects. In comparison, our method not only overcomes the challenging problem of dramatic topology changes in interactive scenes but also maintains the ability to reconstruct high-quality static scenes.
5.2 Language Grounding Comparison
We assess the language grounding performance on the OmniSim dataset using mIOU metric. Figure. 8 suggests that our method obtains the highest mIOU score, with an average of 86.86. In contrast, traditional methods like LERF [24] encounter difficulties in locating objects precisely, with an average mIOU of 21.74. Meanwhile, 2D methods like SAM [26] fail to accurately segment the whole target under specific viewing angles, as objects appear discontinuous in the image. Conversely, our method perceives the completeness of the object and has clear knowledge of its boundaries, demonstrating its advantage in language grounding tasks.
mIOU $\uparrow$
setting SAM [26] LERF [24] Ours
OmniSim #easy 61.58 23.60 86.94
#medium 55.13 19.40 86.32
#challenging 63.86 19.87 90.41
#avg 59.11 21.74 86.86
InterReal #medium 93.27 27.63 84.37
#challenging 91.50 34.39 91.90
#avg 92.82 29.32 86.26
Figure 8: Language Grounding Performance on OmniSim Dataset left): Our method gains the highest mIOU score. right): LiveScene’s grounding exhibits clearer boundaries than other methods.
Figure 9: Rendering and Grounding Performance for #1 and #6. above): Multi-scale factorization greatly boosts the performance of RGB rendering and geometry reconstruction. below): Without view consistency, the model struggles when objects have similar appearances.
Figure 10: Learning process of the probability fields from 0 to 1000 training steps. The model progressively converges to the vicinity of the interactive objects, establishing interactive regions.
5.3 Ablation Study
In this section, we present ablative studies to investigate the effectiveness of each component in LiveScene. We selected 1 scene from the #medium subset, 2 scenes from the #easy subset of OmniSim dataset, and 1 scene each from the #medium and #challenging settings for InterReal. Notably, ground-truth state quantities are only available in OmniSim, not in InterReal. Therefore, we use GT quantities on OmniSim and introduce a learnable variable on InterReal to infer state changes. Figure. 9 reports the rendering quality and grounding performance for #1 and #6.
Effectiveness of components. As illustrated in Table. 4, the multi-scale factorization significantly improves the rendering performance on both datasets, with PSNR on OmniSim increasing from 31.74 to 35.094, shown in #1. Introducing learnable variables for each frame (#2) yields corresponding improvements on InterReal dataset since this latent code can perceive the change in object states. The feature repulsion loss and probability rejection (#3 and #4) together make rendering quality better in InterReal as well as in OmniSim dataset. As for grounding, #5 shows that rendering embeddings along a ray [27, 71] struggles to locate objects precisely. Ensuring view consistency further boosts grounding performance on OmniSim, as demonstrated in #6.
Probability field training. We provide additional experiments in Figure. 10 of the disjoint regions to illustrate the learning process of the probability field from 0 to 1000 training steps. The results demonstrate a clear trend that, as training advances, the proposed method can progressively converge to the vicinity of the interactive objects, thereby establishing interactive regions. With the establishment of the probability field, the model can focus on different interactive objects and guide the sampling process by maximizing probability, thereby achieving disentanglement of interactive scenes.
Table 4: Ablation Study on the subset of InterReal and OmniSim Datasets.
#exp #settings (rendering) InterReal OmniSim
[…]
#0 25.329 0.731 0.329 31.740 0.938 0.118 0.238
#1 28.289 0.819 0.226 35.094 0.969 0.059 0.086
#2 29.577 0.865 0.162
#3 29.959 0.883 0.131 34.989 0.969 0.059 0.085
#4 30.123 0.884 0.132 34.977 0.967 0.061 0.086
LiveScene 30.591 0.896 0.115 35.254 0.971 0.057 0.042
#settings (grounding) mIOU $\uparrow$ mIOU $\uparrow$
#5 30.40 32.87
#6 93.10 71.64
LiveScene 93.02 78.52
[…]
6 Conclusion and Limitation
We present LiveScene, the first language-embedded interactive neural radiance field for complex scenes with multiple interactive objects. A parameter-efficient factorization technique is proposed to decompose interactive spaces into local deformable fields to model individual interactive objects. Moreover, we introduce a novel interaction-aware language embedding mechanism that effectively localizes and controls interactive objects using natural language. Finally, We construct two challenging datasets that contain multiple interactive objects in complex scenes and evaluate the effectiveness and robustness of LiveScene.
Limitations: The control ability of LiveScene is limited by label density. Additionally, our natural language control is currently restricted to closed vocabulary, and it is inherently tied to the capabilities of the underlying foundation model, such as OpenCLIP. In future work, we plan to extend our method to enable open-vocabulary grounding and control, increasing the model’s flexibility and range of applications.
Acknowledgements. This work is partially supported by the Shanghai AI Laboratory, National Key R&D Program of China (JF-P23KK00072-3-DF), the National Natural Science Foundation of China (62376222), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
[…]
Supplementary Material
7 Configurations of OmniSim and InterReal datasets
Figure 11: Illustration of the proposed Omniverse behavior synthetic (OminiSim) and Real captured interactive (InterReal) dataset. These datasets are captured in an OmniGibson simulator or real scene and carefully annotated, providing #28 interactive subsets with 2 Million samples, including RGB, depth, segmentation, camera trajectory, interaction variables, and object captions modalities.
Scene Assets and Generation Pipeline for OmniSim. We generate the synthetic dataset using the OmniGibson simulator. The dataset consists of 20 interactive scenes from 7 scene models: #rs, #ihlen, #beechwood, #merom, #pomaria, #wainscott, and #benevolence. The scenes feature various interactive objects, including cabinets, refrigerators, doors, drawers, and more, each with different hinge joints.
We configure the simulator camera with an intrinsic parameter set of focal length 8, aperture 20, and a resolution of 1024 × 1024. By varying the rotation vectors for each joint of the articulated objects, we can observe different motion states of various objects. We generated 20 high-definition subsets, each consisting of RGB images, depth, camera trajectory, interactive object masks, and corresponding object state quantities relative to their "closed" state at each time step, from multiple camera trajectories and viewpoints.
The data is obtained through the following steps: 1) The scene model is loaded, and the respective objects are selected, with motion trajectories set for each joint. 2) Keyframes are set for camera movement in the scene, and smooth trajectories are obtained through interpolation. 3) The simulator is then initiated, and the information captured by the camera at each moment is recorded.
Scene Assets and Generation Pipeline for InterReal. InterReal is primarily captured using the Polycam app on an Apple iPhone 15 Pro. We selected 8 everyday scenes and placed various interactive objects within each scene, including transformers, laptops, microwaves, and more. We recorded 8 videos, each at a frame rate of 5FPS, capturing 700 to 1000 frames per video.
The dataset was processed via the following steps: 1) manual object movement and keyframe capture, 2) OBJ file export and pose optimization using Polycam, 3) conversion to a dataset containing RGB images and transformation matrices using Nerfstudio [54], and 4) mask generation for each object in each scene using SAM [26] and corresponding prompts and state quantity labeling for certain keyframes.
Statistic of OmniSim and InterReal Datasets. The detailed statistics of the OmniSim and InterReal datasets are shown in Table. 5. The OmniSim dataset consists of 20 interactive scenes, each with 2 to 6 objects, and the InterReal dataset contains 8 real-world scenes, each with 1 to 3 objects. The datasets include RGB, depth, pose, mask, and text prompts modalities, providing a total of 2 million samples for training and evaluation. The objects in the datasets include cabinets, refrigerators, doors, drawers, transformers, laptops, microwaves, and more, with various interactive states and captions.
Table 5: Statistic of OmniSim and InterReal Datasets.
datasets #objects #frame #key frame value rgb depth pose mask text prompts
OmniSim #seq001_Rs_int 4 770 770 fridge, microwave, oven, top cabinet
#seq002_Rs_int 4 2190 2190 fridge, microwave, oven, top cabinet
#seq003_Ihlen_1_int 3 1610 1610 bottom cabinet, dishwasher, top cabinet
#seq004_Ihlen_1_int 2 1630 1630 bottom cabinet, cedar chest
#seq005_Beechwood_0_int 2 1370 1370 bottom cabinet, door
#seq006_Beechwood_0_int 2 1610 1610 dishwasher, microwave
#seq007_Beechwood_0_int 3 1450 1450 bottom cabinet, door, top cabinet
#seq008_Benevolence_1_int 4 1830 1830 door, fridge, microwave, top cabinet
#seq009_Benevolence_1_int 2 1690 1690 cedar chest, door
#seq010_Merom_1_int 3 1930 1930 dishwasher, fridge, microwave, top cabinet
#seq011_Merom_1_int 3 1690 1690 bottom cabinet, top cabinet, door
#seq012_Pomaria_1_int 2 970 970 bottom cabinet, fridge
#seq013_Pomaria_1_int 3 770 770 bottom cabinet, fridge
#seq014_Wainscott_0_int 2 1850 1850 bottom cabinet, cedar chest
#seq015_Wainscott_0_int 2 1350 1350 bottom cabinet, door
#seq016_Wainscott_0_int 2 1170 1170 fridge, stove
#seq017_Benevolence_1_int 6 4590 4590 cedar chest, door, door, fridge, microwave, top cabinet
#seq018_Benevolence_1_int 2 2050 2050 door, top cabinet
#seq019_Rs_int 2 1130 1130 fridge, top cabinet
#seq020_Merom_1_int 2 1990 1990 bottom cabinet, door
#demo 5 6267 6267 cedar chest, door, fridge, oven, top cabinet
#demo001 4 2040 2040 fridge, microwave, oven, top cabinet
#demo002 3 2395 2395 dishwasher, microwave, top cabinet
#demo003 3 2480 2480 dishwasher, oven, top cabinet
#demo004 3 2280 2280 dishwasher, fridge, stove
#demo005 3 1670 1670 bottom cabinet, cedar chest, door
InterReal #seq001_transformer 1 329 38 yellow toy car
#seq002_transformer 1 329 43 blue toy car
#seq003_door 1 355 31 door
#seq004_dog 1 213 41 black mechanical dog
#seq005_sit 1 913 25 small white humanoid
#seq006_stand 1 899 33 small white humanoid
#seq007_flower 3 620 153 blue toy car, yellow toy car, black laptop
#seq008_office 4 1087 658 blue toy car, yellow toy car, black laptop, microwave
8 Videos Demonstration and Anonymous Link
We provide a video of our proposed method LiveScene along with this document to demonstrate the interactive scene reconstruction and multimodal control capabilities. Please refer to the anonymous link: https://livescenes.github.io for more information.
9 Additional implementation details
Loss Functions. In this section, we provide detailed descriptions of the loss functions used in LiveScene:
$\mathcal{L}_{total}=\mathcal{L}_{\text{MSE}}+\lambda_{1}\mathcal{L}_{\text{focus}}+\lambda_{2}\mathcal{L}_{\text{repuls}}+\lambda_{3}\mathcal{L}_{\text{var}}+\lambda_{4}\mathcal{L}_{\text{lang}}+\lambda_{5}\mathcal{L}_{\text{smooth}},$ (6)
Rendering Loss. We use the standard NeRF rendering loss, which is the sum of the mean squared error (MSE) between the rendered color and the ground truth color, and the MSE between the rendered depth and the ground truth depth. The loss is computed for each pixel in the image and averaged over the entire image:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{C}_{i}-\tilde{\mathbf{C}}_{i}\right\|^{2},$ (7)
where $\mathbf{C}_{i}$ and $\tilde{\mathbf{C}}_{i}$ are the rendered and ground truth RGB values, respectively, and $N$ is the number of pixels in the image.
Focal Loss. Due to the predominance of background regions in the images, we employ focal loss to enhance the model’s focus on the relatively smaller interactive mask regions:
$\mathcal{L}_{\text{focus}}=\beta\cdot\left(1-e^{\sum_{i=1}^{\alpha}\mathbf{M}_{i}\log(\hat{\mathbf{P}}_{i})}\right)^{\gamma}\cdot\left(-\sum_{i=1}^{\alpha}\mathbf{M}_{i}\log(\hat{\mathbf{P}}_{i})\right),$ (8)
where $\mathbf{M}$ is the ground truth mask label, $\hat{\mathbf{P}}$ is the probability map rendering from the interactive probability field, $\beta$ is the balancing factor, and $\gamma$ is the focusing parameter. In our experiments, we set $\alpha=0.5$ and $\gamma=1.5$.
Repulsion Loss. To avoid sampling conflicts and feature oscillations at the boundaries, we introduce a repulsion loss to amplify the feature differences between distinct deformable scenes, thereby promoting the separation of deformable field:
$\mathcal{L}_{\text{repuls}}=\mathbf{ELU}(K-\left\|(\mathbf{M}_{i}\odot\mathbf{M}_{j})(\mathcal{F}_{i}-\mathcal{F}_{j})\right\|),\end{aligned}$
where $\mathbf{M}_{i}$ and $\mathbf{M}_{j}$ are the ground truth mask of rays, and $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ are the last-layer features of interaction probability decoder in Figure. 2. $K$ is the constant hyperparameters. In training iteration, we randomly select ray pairs and apply $\mathcal{L}_{\text{repuls}}$ to enforce the separation of interactive probability features across local deformable spaces. Our approach draws inspiration from [25], which has shown the effectiveness of repulsive forces in resolving ambiguities in 3D segmentation.
Interaction Variable MSE. We follow the value MSE in [21] and use the standard MSE loss to supervise the interaction values training:
$\mathcal{L}_{\text{var}}=\frac{1}{N}\sum_{i=1}^{N}\left\|\boldsymbol{\kappa}_{i}-\tilde{\boldsymbol{\kappa}}_{i}\right\|^{2},$ (10)
where $\boldsymbol{\kappa}_{i}$ and $\tilde{\boldsymbol{\kappa}}$ are the predicted and ground truth interaction variables, respectively, and $N$ is the number of ray samples of a batch. Note that we only apply $\mathcal{L}_{\text{var}}$ to the InterReal dataset and use learnable variables as inputs to the model due to the lack of dense ground truth interaction variables. In OmniSim, we directly use the ground truth interaction variables as inputs provided by the simulator to achieve precise control.
Language Embedding L2 Loss. In LiveScene implementation, we use the huber loss to supervise the language embedding training. But we do not distill the language embedding in the 3D language field but we store the language embedding in the proposed interaction-aware language feature plane. The loss is defined as:
$\mathcal{L}_{\text{lang}}(\phi(\mathbf{p}),\tilde{\phi}(\mathbf{p}))=\left\{\begin{array}[]{ll}\frac{1}{2}(\phi(\mathbf{p})-\phi(\mathbf{p}))^{2}&\text{ if }|\phi(\mathbf{p})-\tilde{\phi}(\mathbf{p})|\leq\delta\\ \delta\left(|\phi(\mathbf{p})-\tilde{\phi}(\mathbf{p})|-\frac{1}{2}\delta\right)&\text{ otherwise }\end{array}\right.,$ (11)
where $\phi(\mathbf{p})$ and $\tilde{\phi}(\mathbf{p})$ are the predicted and ground truth language embeddings, respectively, and $\delta$ is the threshold. In our experiments, we set $\delta=1.0$.
Smoothness Loss. Inspired by K-Planes [9], we use 1D Laplacian (second derivative) filter to smooth the local deformable field feature plane, which helps to reduce the noise in the deformable field and alleviate feature oscillations and sampling conflicts at the sampling boundary:
$\mathcal{L}_{\text{smooth}}(\mathbf{p})=\frac{1}{|L|n^{2}}\sum_{l,i,k}\left\|\mathbf{p}_{l}^{i,k-1}-2\mathbf{p}_{l}^{i,k}+\mathbf{p}_{l}^{i,k+1}\right\|_{2}^{2},$ (12)
where $i$ and $k$ are indices on the plane resolution $n$, and $l$ is the feature planes index.
Probability Rejection Operation. Additionally, a probability rejection operation is proposed to truncate the low-probability samples if the deformable probability at $\mathbf{p}$ is smaller than threshold $s$. The probability rejection is proposed to truncate the low-probability samples if the maximum deformable probability $\mathbf{P}$ at $\mathbf{p}$ is smaller than threshold $s$ and selects the background feature directly. The operation is defined as:
$u=\left\{\begin{array}[]{ll}\operatorname*{arg\,max}_{i}\{\mathbf{P}_{i}\},&\text{if}\quad\mathbf{P}_{i}\geq s\\ -1&\text{ otherwise }\end{array}\right..$ (13)
Implementation Details. LiveScene is implemented in Nerfstudio [54] from scratch. We represent the field as a multi-scale feature plane with resolutions of $512\times 256\times 128$, and feature dimension of 32. The proposal network adopts a coarse-to-fine sampling process, where each sampling step concatenates the position feature and the state quantity as the query for the 4D deformation mask field, which is a 1-layer MLP with 64 neurons and ReLU activation. For InterReal, we introduce additional learnable variables bound to each frame to capture changes in object states within the scene. These variables are represented by a plane with a resolution typically half the frame number, with a feature dimension of 4 for most scenes. For all experiments, we use the Adam optimizer with initial learning rates of 0.01 and a cosine decay scheduler with 512 warmup steps for all networks. We set loss weights as follows: $\lambda_{1}=1e-3,\lambda_{2}=1e-2,\lambda_{3}=1e-3,\lambda_{4}=1.0,\lambda_{5}=1e-3$. The model is trained for 80k steps on the OmniSim dataset and 100k steps on the InterReal dataset, using a batch size of 4096 rays with 64 samples each. We run the model on an NVIDIA A100 GPU, requiring approximately 4 hours and 40GB of memory.
10 Additional Experimental Results
Model Parameter Efficiency Comparison. We compare the number of parameters of LiveScene with other methods in Table. 6, varying the number of interactive objects in the scene. The results show that LiveScene has a constant number of parameters, regardless of the number of interactive objects, making it more efficient than other methods. In contrast, CoGS [65] has a higher base number of parameters and a linear increase with the number of interactive objects. MK-Planes [9] exhibits a quadratic increase in parameters with the number of interactive objects. Although NeRF[40] and InstantNGP [41] have a low number of parameters, they are limited to 3D static scene reconstruction.
Table 6: Model Parameters vs Object Quantity.
Method # interactive objects Trend
1 2 3 4 5 6
NeRF [40] 13.23 13.23 13.23 13.23 13.23 13.23 Constant
InstantNGP [41] 11.68 11.68 11.68 11.68 11.68 11.68 Constant
K-Planes [9] 35.66 35.66 35.66 35.66 35.66 35.66 Constant
MK-Planes 35.66 35.96 36.32 36.76 37.27 37.86 Quadratic
MK-Planes 35.66 35.88 36.10 36.32 36.54 36.76 Linear
CoGS [65] 42.24 43.23 44.23 45.23 46.23 47.22 Linear
LiveScene (Ours) 34.52 34.52 34.52 34.52 34.52 34.52 Constant
View Synthesis Quality Comparison on OmniSim and InterReal dataset We provide detailed quantitative results on the OmniSim and InterReal datasets in Table. 7 and Table. 8 provide detailed quantitative results on the OmniSim and InterReal datasets, respectively. LiveScene outperforms prior works on most metrics and achieves the best PSNR on the #challenging and #office subsets with a significant margin. Note that the #challenging and #office subsets contain scenes with multiple interactive objects and large deformable fields, which are challenging for existing methods. We report the score as NaN if the model fails to converge or is out of memory during training multiple times.
Table 7: Detailed Quantitative results on OmniSim Dataset. LiveScene outperforms prior works on most metrics and achieves the best PSNR on the #challenging subset with a significant margin.
Dataset Metric NeRF [40] Instant-NGP [41] HyperNeRF [44] K-Planes [9] CoNeRF [21] MK-Planes [9] MK-Planes [9] CoGS [65] LiveScene
#seq001_Rs PSNR 25.941 25.768 NaN 33.136 34.035 32.169 32.092 32.211 34.784
#seq001_Rs SSIM 0.931 0.933 NaN 0.953 0.957 0.946 0.946 0.968 0.974
#seq001_Rs LPIPS 0.118 0.113 NaN 0.093 0.135 0.110 0.110 0.068 0.048
#seq002_Rs PSNR 28.616 28.660 NaN 34.765 34.286 36.532 34.580 34.497 35.190
#seq002_Rs SSIM 0.950 0.946 NaN 0.967 0.951 0.976 0.968 0.979 0.969
#seq002_Rs LPIPS 0.096 0.112 NaN 0.074 0.217 0.036 0.074 0.051 0.070
#seq003_Ihlen PSNR 26.720 28.255 33.551 35.217 34.700 34.758 34.753 36.816 35.323
#seq003_Ihlen SSIM 0.940 0.944 0.946 0.964 0.953 0.966 0.966 0.980 0.966
#seq003_Ihlen LPIPS 0.120 0.121 0.268 0.097 0.244 0.087 0.090 0.077 0.094
#seq004_Ihlen PSNR 30.847 31.800 31.115 36.157 32.684 34.863 35.000 31.055 36.712
#seq004_Ihlen SSIM 0.927 0.942 0.878 0.955 0.888 0.919 0.926 0.915 0.962
#seq004_Ihlen LPIPS 0.104 0.102 0.389 0.085 0.366 0.145 0.135 0.209 0.072
#seq005_Beechwood PSNR 27.183 27.295 30.699 31.944 32.549 33.195 33.098 33.664 33.623
#seq005_Beechwood SSIM 0.930 0.937 0.906 0.944 0.927 0.961 0.959 0.978 0.962
#seq005_Beechwood LPIPS 0.127 0.112 0.291 0.105 0.245 0.076 0.080 0.058 0.072
#seq006_Beechwood PSNR 27.988 28.150 29.513 31.861 30.058 31.541 31.521 31.272 32.206
#seq006_Beechwood SSIM 0.938 0.938 0.907 0.951 0.917 0.951 0.951 0.974 0.959
#seq006_Beechwood LPIPS 0.103 0.119 0.314 0.097 0.283 0.095 0.096 0.059 0.077
#seq007_Beechwood PSNR 23.201 22.902 31.259 30.979 33.451 30.136 30.089 27.367 30.360
#seq007_Beechwood SSIM 0.885 0.886 0.913 0.938 0.935 0.942 0.942 0.893 0.946
#seq007_Beechwood LPIPS 0.220 0.219 0.289 0.140 0.229 0.120 0.121 0.219 0.107
#seq008_Benevolence PSNR 25.750 25.574 32.691 31.914 34.319 30.926 30.916 33.795 33.393
#seq008_Benevolence SSIM 0.943 0.940 0.945 0.948 0.960 0.941 0.941 0.980 0.970
#seq008_Benevolence LPIPS 0.113 0.123 0.229 0.107 0.185 0.118 0.116 0.072 0.067
#seq009_Benevolence PSNR 24.326 24.386 29.596 32.836 31.225 31.500 31.471 33.205 32.030
#seq009_Benevolence SSIM 0.921 0.922 0.897 0.956 0.932 0.954 0.953 0.975 0.962
#seq009_Benevolence LPIPS 0.124 0.128 0.327 0.090 0.248 0.088 0.090 0.074 0.071
#seq010_Merom PSNR 22.927 22.765 28.985 30.120 31.092 29.461 29.396 30.254 30.029
#seq010_Merom SSIM 0.917 0.925 0.939 0.960 0.957 0.960 0.959 0.974 0.966
#seq010_Merom LPIPS 0.173 0.158 0.275 0.093 0.233 0.087 0.088 0.065 0.074
#seq011_Merom PSNR 26.732 27.077 NaN 33.394 30.483 32.951 32.910 31.767 33.426
#seq011_Merom SSIM 0.932 0.933 NaN 0.959 0.932 0.959 0.959 0.968 0.960
#seq011_Merom LPIPS 0.112 0.117 NaN 0.074 0.246 0.073 0.072 0.091 0.068
#seq012_Pomaria PSNR 26.856 27.074 NaN 35.185 33.065 32.248 32.209 37.284 33.367
#seq012_Pomaria SSIM 0.936 0.943 NaN 0.972 0.954 0.966 0.966 0.985 0.969
#seq012_Pomaria LPIPS 0.138 0.126 NaN 0.059 0.199 0.075 0.075 0.047 0.061
#seq013_Pomaria PSNR 25.277 24.018 NaN 30.860 33.682 30.390 30.299 32.868 33.592
#seq013_Pomaria SSIM 0.925 0.930 NaN 0.943 0.964 0.931 0.930 0.981 0.970
#seq013_Pomaria LPIPS 0.154 0.161 NaN 0.123 0.166 0.162 0.164 0.045 0.056
#seq014_Wainscott PSNR 26.011 25.966 NaN 32.517 29.580 30.511 30.504 31.885 31.197
#seq014_Wainscott SSIM 0.927 0.924 NaN 0.955 0.925 0.951 0.951 0.969 0.952
#seq014_Wainscott LPIPS 0.105 0.116 NaN 0.077 0.244 0.082 0.083 0.067 0.083
#seq015_Wainscott PSNR 27.257 27.191 NaN 30.721 32.307 28.288 28.134 32.949 34.266
#seq015_Wainscott SSIM 0.953 0.951 NaN 0.955 0.962 0.942 0.942 0.975 0.976
#seq015_Wainscott LPIPS 0.080 0.092 NaN 0.083 0.202 0.110 0.108 0.078 0.050
#seq016_Wainscott PSNR 21.953 21.660 28.364 30.414 30.205 28.915 28.710 31.965 29.746
#seq016_Wainscott SSIM 0.897 0.895 0.909 0.951 0.935 0.952 0.951 0.976 0.955
#seq016_Wainscott LPIPS 0.175 0.194 0.327 0.089 0.260 0.086 0.087 0.066 0.083
#seq017_Benevolence PSNR 26.364 26.367 27.533 29.833 30.349 29.254 26.565 28.701 31.645
#seq017_Benevolence SSIM 0.927 0.920 0.897 0.937 0.923 0.933 0.887 0.970 0.948
#seq017_Benevolence LPIPS 0.128 0.143 0.318 0.118 0.238 0.119 0.218 0.073 0.093
#seq018_Benevolence PSNR 28.236 24.296 32.551 34.690 34.297 33.049 33.002 34.963 34.187
#seq018_Benevolence SSIM 0.918 0.809 0.911 0.951 0.936 0.953 0.952 0.976 0.958
#seq018_Benevolence LPIPS 0.145 0.342 0.293 0.093 0.248 0.090 0.091 0.114 0.081
#seq019_Rs PSNR 20.059 20.854 33.119 34.462 34.598 33.679 33.653 25.947 35.223
#seq019_Rs SSIM 0.794 0.808 0.950 0.956 0.963 0.963 0.962 0.879 0.969
#seq019_Rs LPIPS 0.425 0.424 0.270 0.106 0.225 0.087 0.089 0.327 0.068
#seq020_Merom PSNR 23.273 24.074 31.280 30.462 32.580 30.655 30.626 31.280 32.869
#seq020_Merom SSIM 0.823 0.852 0.970 0.929 0.914 0.919 0.918 0.970 0.954
#seq020_Merom LPIPS 0.306 0.259 0.086 0.140 0.276 0.139 0.142 0.086 0.078
Table 8: Detailed Quantitative results on InterReal Dataset. Our method outperforms others in most settings, with a significant advantage of PSNR, SSIM, and LPIPS on the #challenging subset.
dataset Metric NeRF [40] Instant-NGP [41] HyperNeR [44]F K-Planes [9] CoNeRF [21] CoGS [21] LiveScene
#seq01_transformer PSNR 20.094 20.619 24.651 26.881 27.260 31.067 30.396
#seq01_transformer SSIM 0.725 0.805 0.638 0.791 0.739 0.943 0.912
#seq01_transformer LPIPS 0.182 0.167 0.495 0.185 0.355 0.060 0.060
#seq02_transformer PSNR 20.093 20.028 24.433 26.232 26.917 30.513 29.706
#seq02_transformer SSIM 0.736 0.778 0.635 0.763 0.732 0.938 0.899
#seq02_transformer LPIPS 0.210 0.196 0.477 0.223 0.357 0.062 0.069
#seq03_door PSNR 20.001 20.652 27.144 29.278 29.850 31.998 32.709
#seq03_door SSIM 0.785 0.831 0.878 0.920 0.922 0.962 0.960
#seq03_door LPIPS 0.250 0.250 0.316 0.101 0.231 0.071 0.044
#seq04_dog PSNR 20.044 20.206 25.691 30.350 28.567 32.455 32.519
#seq04_dog SSIM 0.723 0.819 0.730 0.894 0.815 0.950 0.943
#seq04_dog LPIPS 0.196 0.178 0.435 0.107 0.324 0.074 0.049
#seq05_sit PSNR 21.558 24.211 24.944 27.970 26.252 27.169 30.161
#seq05_sit SSIM 0.480 0.727 0.573 0.773 0.633 0.767 0.886
#seq05_sit LPIPS 0.178 0.236 0.543 0.207 0.463 0.232 0.084
#seq06_stand PSNR 23.109 24.483 24.833 27.285 26.159 31.442 29.400
#seq06_stand SSIM 0.643 0.699 0.574 0.736 0.627 0.919 0.868
#seq06_stand LPIPS 0.123 0.260 0.538 0.237 0.470 0.104 0.089
#seq07_flower PSNR 21.150 21.813 25.334 26.545 26.854 28.435 28.208
#seq07_flower SSIM 0.721 0.747 0.712 0.759 0.748 0.893 0.844
#seq07_flower LPIPS 0.302 0.319 0.489 0.321 0.425 0.165 0.188
#seq08_office PSNR 21.187 21.474 25.188 26.309 26.040 27.510 28.663
#seq08_office SSIM 0.735 0.743 0.714 0.754 0.720 0.897 0.848
#seq08_office LPIPS 0.371 0.358 0.545 0.341 0.520 0.138 0.181
Interactive Scenes Geometry Comparison . To evaluate the completeness of the topological structure of interactive objects, we employ the depth L1 error metric. As shown in Figure. 12, our method outperforms SOTA methods on scenes from OmniSim. While existing methods excel in RGB image rendering, they struggle with depth structure representation. Specifically, CoNeRF [21] performs relatively well in #seq08 and #seq15 but fails in large-scale scenes (#seq17 and #seq19). CoGS [65] exhibit notable artifacts in the depth map. Moreover, MK-Planes also fails to recover depth around interactive objects. In contrast, our method achieves the lowest depth error and renders satisfying depth maps, demonstrating its accurate interactive scene modeling capabilities.
Language Grounding Comparison . We assess the language grounding performance on OmniSim dataset using mIOU metric. Figure. 13 suggests that our method obtains the highest mIOU score, with an average of 86.86. In contrast, traditional methods like LERF [24] encounter difficulties in locating objects precisely, with an average mIOU of 21.74. Meanwhile, 2D methods like SAM [26] fail to accurately segment the whole target under specific viewing angles, as objects appear discontinuous in the image. Conversely, our method perceives the completeness of the object and has clear knowledge of its boundaries, demonstrating its advantage in language grounding tasks.
Depth L1 Error $\downarrow$
scene CoNeRF MK-Planes CoGS LiveScene
#seq04 0.029 0.037 0.381 0.018
#seq08 0.018 0.301 0.655 0.025
#seq14 0.042 0.103 0.810 0.039
#seq15 0.019 0.568 0.690 0.021
#seq17 0.282 0.706 0.019
#seq19 0.136 0.689 0.034
#avg 0.027 0.238 0.655 0.026
Figure 12: Structure Reconstruction Performance on OmniSim Dataset. Our method surpasses most previous works on chosen subsets.
mIOU $\uparrow$
setting SAM [26] LERF [24] Ours
OmniSim #easy 61.58 23.60 86.94
#medium 55.13 19.40 86.32
#challenging 63.86 19.87 90.41
#avg 59.11 21.74 86.86
InterReal #medium 93.27 27.63 84.37
#challenging 91.50 34.39 91.90
#avg 92.82 29.32 86.26
Figure 13: Language Grounding Performance on OmniSim Dataset. left): Our method gains the highest mIOU score. right): LiveScene’s grounding exhibits clearer boundaries than other methods.
More Detailed Rendering Comparison We provide more detailed visual comparisons, including RGB, depth, and language grounding on the OmniSim and InterReal datasets in Figure. 14, Figure. 15, Figure. 16, and Figure. 17, respectively. Our method surpasses existing approaches by reconstructing more detailed and accurate representations of the objects. In both datasets, LiveScene can generate more accurate and detailed object shapes and textures, especially for scenes with multiple interactive objects and large deformable fields. Compared with LERF [24], our method can generate more accurate language grounding results, which is crucial for interactive object manipulation tasks, demonstrated in Figure. 17.
Figure 14: View Synthesis Visualization on InterReal Dataset. We compare our method with SOTA methods on RGB rendering across real scenes. LiveScene obtained more detailed and accurate representations of the objects. While other methods fail to capture the object’s shape and cause significant artifacts.
Figure 15: View Synthesis Visualization on InterReal Dataset. compared with the other methods, LiveScene reconstructs clear and accurate object shapes and textures.
Figure 16: Illustration of the depth map comparison on the OmniSim datasets. Our method can generate more accurate depth maps than other methods, demonstrating the effectiveness of interactive scene reconstruction. In contrast, other methods either fail to capture the object’s shape or cause significant artifacts.
Figure 17: Illustration of the language grounding comparison on the OmniSim datasets. Compared to LeRF, our method can locate more accurate interactive objects, overcoming the obvious inconsistency problem in interactions, while maintaining accurate boundaries. In contrast, LeRF suffers from a diffusion phenomenon in object localization due to changes in object topology structure.
10.1 More Ablation Studies
4D Deformable Feature Visualization. We provide additional interaction feature visualization of x-z, y-z, and z-k in Figure. 18(a) to illustrate latent feature distribution. It can be seen that the features are clustered around the spatial coordinates of interactive objects, corresponding to the local deformable fields in Sec 3.2 of the manuscript. Figure. 18(b) validates the performance of LiveScene in scenarios with up to 10 complex interactive objects. Notably, our method demonstrates robustness in rendering quality, which does not degrade significantly as the object number increases. The number of objects is not a major limiting and our method is still feasible as long as the dataset provides mask and control variable labels. In contrast, the occlusion and topological complexity between objects do affect the reconstruction results, which will be discussed in the limitations section. In Figure. 18(c), we demonstrate the fine-grained control capability of LiveScene on a refrigerator and cabinet dataset without part-based labels. Our method can control a part of the object even though there are no individual part-based interaction variable labels. However, the effect is not entirely satisfactory, due to the lack of labels and CLIP’s limited understanding of spatial relationships.
Ablation Study on Multi-scale Factorization. We conduct more ablation studies on the OmniSim dataset to evaluate the effectiveness of the multi-scale factorization. The results show that the multi-scale factorization can improve the model’s performance by capturing the object’s detailed structure and texture. However, the model without multi-scale factorization performs poorly in depth rendering, illustrating the improvements of multi-scale factorization in scene geometric modeling. The results are shown in Figure. 19.
Ablation Study on Interaction-aware Language Embedding. We conduct more ablation studies on the OmniSim dataset to evaluate the effectiveness of the interaction-aware language embedding. The results show that the interaction-aware language embedding can effectively improve the model’s performance when encouraging significant scene topological changes. While the model without interaction-aware language embedding fails to ground the correct object because of the lack of interaction-aware information. The results are shown in Figure. 20.
Maximum Probability Embeds Retrieval. We conduct more ablation studies on the OmniSim dataset to evaluate the effectiveness of the maximum probability embedding retrieval. The results show that the maximum probability embedding retrieval can improve the model’s performance with higher storage efficiency and training speed, and the grounding results will also be more concentrated in the object region. The fundamental reason is that this method decouples language from the 3D scene to the object level, rather than the entire 3D space. The results are shown in Figure. 21.
Figure 18: a) Visualization of x-$\kappa$, y-$\kappa$ and z-$\kappa$ latent feature planes. b) Sce ne with more objects. c) Part-level interaction.
Figure 19: More ablation of multi-scale factorization on the OmniSim dataset. We compare the performance of LiveScene with w/o multiscale factor. The results show that the multi-scale factorization can improve the model’s performance by capturing the object’s detailed structure and texture. However, the model without multi-scale factorization performs poorly in depth rendering, illustrating the improvements of multi-scale factorization in scene geometric modeling.
Figure 20: More ablation of interactive object modeling on the OmniSim dataset. We compare the performance of LiveScene with w/o interaction-aware language embedding. The results show that the interaction-aware language embedding can effectively improve the model’s performance when encouraging significant scene topological changes. While the model without interaction-aware language embedding fails to ground the correct object because of the lack of interaction-aware information.
Figure 21: By applying the proposed multiscale factor and maximum probability embedding retrieval, the model achieves better performance with higher storage efficiency and training speed, and the grounding results will also be more concentrated in the object region. The fundamental reason is that this method decouples language from the 3D scene to the object level, rather than the entire 3D space.